OpenStudy (anonymous):

Use the properties of exponents to simplify the expression. 10^x+1 divide by 10^x-3

OpenStudy (anonymous):

The divisional property of exponents is that when you divide by an exponent (when the base is equal) that you subtract the exponent being divided.

OpenStudy (anonymous):

Wait, I'm an idiot, lol.

OpenStudy (anonymous):

yeah i get that but i am not sure what to do when it is a equation

OpenStudy (anonymous):

Same thing, you just substitute the whole equation.

OpenStudy (anonymous):

If x = 1, what you would have is 10^2 / 10^-2, right?

OpenStudy (anonymous):

Then you would have 10^2-(-2)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Since we don't x, we leave it there.

OpenStudy (anonymous):

Since we don't *know* x

OpenStudy (anonymous):

So if it was 10^A / 10^B, we did 10^A-B

OpenStudy (anonymous):

But A was "X + 1" and B was "X - 3"

OpenStudy (anonymous):

oh i didnt distribute the negative

OpenStudy (anonymous):

So we have 10^x+1 / 10^x-3, becomes 10^(x+1)-(x-3), -(x-3) = -x + 3, right? So we have 10^x+1-x+3 Finally: We have 10 to all these together: x -x +1 +3 x + (-x) = 0 1 + 3 = 4 0+4 = 4 So we have 10^4

OpenStudy (anonymous):

Sorry, had to correct a number typo :)

OpenStudy (anonymous):

Yea, that's easy to do, haha. Everything figured out?

OpenStudy (anonymous):

yes... thank you

OpenStudy (anonymous):

Awesome :)